# Aggregation Artifacts in Subjective Tasks Collapse Large Language Models’ Posteriors

### v1

@@ Abstract
In-context Learning (ICL) has become the primary method for performing natural language tasks with Large Language Models (LLMs). The knowledge acquired during pre-training is crucial for this few-shot capability, providing the model with task priors. However, recent studies have shown that ICL predominantly relies on retrieving task priors rather than “learning” to perform tasks. This limitation is particularly evident in complex subjective domains such as emotion and morality, where priors significantly influence posterior predictions. In this work, we examine whether this is the result of the aggregation used in corresponding datasets, where trying to combine low-agreement, disparate annotations might lead to annotation artifacts that create detrimental noise in the prompt. Moreover, we evaluate the posterior bias towards certain annotators by grounding our study in appropriate, quantitative measures of LLM priors. Our results indicate that aggregation is a confounding factor in the modeling of subjective tasks, and advocate focusing on modeling individuals instead. However, aggregation does not explain the entire gap between ICL and the state of the art, meaning other factors in such tasks also account for the observed phenomena. Finally, by rigorously studying annotator-level labels, we find that it is possible for minority annotators to both better align with LLMs and have their perspectives further amplified.
@@ 1 Introduction
While ICL is often contrasted with traditional in-weights learning (i.e., gradient-based updates of the models’ parameters) Kossen et al. (2023); Chan et al. (2022), the ICL abilities of LLMs depend on their general, in-weights prior knowledge of several concepts, allowing them to perform many tasks in a zero-shot or few-shot manner. Therefore, studying how LLMs leverage the context in relation with their existing knowledge is a prerequisite to understanding ICL.
Prior work found evidence that LLMs may be overly reliant on their prior knowledge, disregarding the demonstrations in the prompt. Specifically, Min et al. (2022) demonstrated that, under certain circumstances, LLMs ignore the provided signal in their prompt in the form of the mapping between inputs and outputs, and instead act as a database of tasks; they focus on the examples and the labels independently to fetch the underlying task Xie et al. (2021): Min et al. (2022) sampled examples and labels independently, and showed very little change in performance. Since no annotations are provided, the setting is given the status of “zero-shot” inference, namely task-recognition zero-shot. While follow-up work Kossen et al. (2023); Wei et al. (2023); Pan et al. (2023); Yoo et al. (2022) has further studied this phenomenon and questioned the generality of the phenomenon, more recent work Chochlakis et al. (2024b, a) has provided further, quantitative evidence for the relative importance of prior and evidence in the posteriors of the models. Specifically, in complex subjective tasks like multilabel emotion or morality recognition, LLMs seem to virtually disregard evidence from the dataset’s mapping in their posterior predictions, even in the form of Chain-of-Thought prompting (CoT; Wei et al. 2022), performing significantly worse than traditional algorithms Alhuzali and Ananiadou (2021). This has implications for annotator modeling, as the same document can receive different but valid annotations by different people.
@@
In this work, we question whether we can efficiently use this crowd truth when modeling these tasks with LLMs. We hypothesize that the aggregation process creates artifacts in the labels provided to the model, and in turn in the prompt, causing the model to ignore the entire input-label mapping as noise. Specifically, aggregation can create inconsistent annotations, since, for instance, different annotators can prevail in different examples, which can cause inconsistencies within or across splits. We create a carefully crafted experimental setting to test our hypothesis with annotator-level labels using ICL and CoT, and try to gauge at the magnitude of this effect. We find that LLMs do indeed tend to consistently favor individual perspectives compared to the aggregate, and in fact favor minority annotators more than majority annotators, who better resemble the aggregate. Our contributions are as follows:
@@ 2.1 In-Context Learning and Priors
ICL Brown et al. (2020) has been used extensively to evaluate LLMs on standard benchmarks Srivastava et al. (2022). It requires no gradient-based interventions, which are otherwise costly to perform for large models, and usually achieves competitive or state-of-the-art performance. The existence of commercial APIs Achiam et al. (2023), coupled with open-weights alternatives Touvron et al. (2023); Dubey et al. (2024) have made ICL an accessible generalist for language tasks and more Liu et al. (2024). Previous work has further studied controlling the reliance on context and in-weights knowledge through the distribution of the training data Chan et al. (2022), examining how to optimally select examples for the prompt Rubin et al. (2022); Gupta et al. (2023), integrating instructions explicitly during training Touvron et al. (2023); Ouyang et al. (2022b), etc. Relevant to our work, prior work has elicited ICL priors by providing random labels for the examples of the prompt Min et al. (2022). The resulting minimal variations in performance indicates that LLMs recognize and retrieve their prior knowledge of the task in the prompt rather than doing any “learning”. Subsequent results challenged the view that LLMs mostly perform task recognition, showcasing a significant degradation in performance when scaling the prompt Kossen et al. (2023) or when focusing on specific tasks instead of aggregates Yoo et al. (2022), and analyzed behavior with certain label manipulations Kossen et al. (2023); Pan et al. (2023); Wei et al. (2023). More recent work, however, suggests that in complex subjective tasks like emotion and morality classification, the prior understanding of the task dominates posterior predictions Chochlakis et al. (2024b).
One potential way to augment ICL and overcome the prior bias is with CoT Wei et al. (2022). CoT incorporates the derivation process from input to output explicitly in the prompt, presenting a more human-like reasoning process. This has several advantages, such as making some patterns in the data explicit in the prompt, making model responses more explainable, and potentially directing more computing resources towards more complex problems. However, detailed analysis has cast some doubt on the reliability and the faithfulness of this reasoning technique Lanham et al. (2023); Turpin et al. (2024). Nonetheless, CoT seems to improve the robustness and performance of LLMs across a plethora of tasks, yet follow-up work on subjective tasks showed no improvements and the same bias towards a reasoning prior, especially for larger models Chochlakis et al. (2024a). Methods such as Tree of Thoughts Yao et al. (2024) or self-consistency Wang et al. (2022) have experimented with ways to further augment CoT. In this work, we study whether part of the prior bias in both ICL and CoT can be explained by the aggregation process used to derive the labels, and in turn the reasoning process, of the examined benchmarks.
@@ 2.2 Annotator Disagreement and Modeling
Many works have attempted to model individual annotator perspectives. Researchers used the EM algorithm Dawid and Skene (1979) to assign confidence to each annotator’s evaluations Hovy et al. (2013). Recently, Gordon et al. (2022) concatenated features derived from Transformers Vaswani et al. (2017) with annotator embeddings that incorporate demographics to model individual perspectives. Demographic information has also been incorporated in word embeddings by  Garten et al. (2019). Demographic information and psychological profiles have also been statistically examined in text annotations to derive insights into systematic biases Sap et al. (2022). Recent work has tried to filter annotators based on deep learning methods Mokhberian et al. (2022), to model annotators on top of common representations Davani et al. (2022); Mokhberian et al. (2023), and to decrease annotation costs based on agreement Golazizian et al. (2024). Modeling annotators with LLMs has shown limited success, and LLM biases have also been explored Dutta et al. (2023); Abdurahman et al. (2024); Hartmann et al. (2023); Tao et al. (2024).
@@ 3 Methodology
We closely follow the methodology and notation of Chochlakis et al. (2024a). For a set of examples $\mathcal{X}$, and a set of labels $\mathcal{Y}$, a dataset $\mathcal{D}^{a}$ defines a mapping $f^{a}:\mathcal{X}\rightarrow\mathcal{Y}$, where $a$ denotes a specific annotator or the aggregate, as well as reasoning chains $R^{a}(x)=r,x\in\mathcal{X}$ that explicitly describe $f^{a}$, and therefore $\mathcal{D}^{a}=\{(x,y,r):x\in\mathcal{X},y=f^{a}(x),r=R^{a}(x)\}$, from which we can sample demonstrations with $p(x,y,r)$. We do not differentiate between splits for brevity. Given CoT prompt $S=\{(x_{i},y_{i},r_{i}):(x_{i},y_{i},r_{i})\sim p_{S},i\in[k]\}$ with $k$ demonstrations sampled with distribution $p_{S}$ from $\mathcal{D}^{a}$ without replacement, an LLM produces its own mapping and predictions for the task, denoted as $\hat{f}_{k}(.;p_{S}):\mathcal{X}\rightarrow\mathcal{Y}$. When using regular ICL, we simply drop the reasoning text. For all our experiments, we set the temperature $\tau=0$ to derive deterministic predictions, and vary the examples and/or labels in the prompt to estimate the variance.
@@ 3.1 Similarity and Performance Metrics
To keep evaluations consistent when using API-based LLMs, we rely on similarity measures calculated directly on the final predictions rather than probabilistic measures like logits. Therefore, we use the Jaccard Score, Micro and Macro F1 metrics Mohammad et al. (2018) to evaluate the performance of the models. For consistency, we also use them to quantify the similarity between the predictions from different model runs or annotators, since they are symmetric functions which we apply to interchangeable sets.
Figure 1: Performance comparison between LLMs w/ and w/o Chain-of-Thought prompting compared to SpanEmo (BERT-based) using aggregated labels.
Figure 2: Scatter plot of annotator similarity with aggregate and with prior.
@@ 3.2 Task Prior Knowledge Proxies via Zero-shot Inference
Here, we precisely define the priors that we use for ICL and CoT. First, we have the true reasoning task-recognition zero-shot111since no annotations are required prior, where the prompt contains $k$ demonstrations sampled with $p^{I}(x,y,r)=p(x)p(y)p(r)$, so text, labels, and reasoning are sampled independently from each other from $\mathcal{D}^{a}$, hence labels and reasoning are irrelevant to the text and each other. This effectively maintains the higher-order relationships between labels, which are strong in such multilabel tasks Cowen and Keltner (2017). For ICL, we have the corresponding task-recognition zero-shot prior sampled with $p^{I}(x,y)=p(x)p(y)$, so text and labels are also sampled independently. The similarity of the priors to annotators and aggregate are, therefore, measured by comparing (as described in Section 3.1) the prior predictions $\hat{f}_{k}(.;p_{I})$ with the annotator’s labels, which is equivalent to the prior performance for the annotator, and the posteriors are simply defined as the $\hat{f}_{k}(.;p_{S})$ with the joint sampling distribution $p_{S}$ (meaning we present the gold labels to the model).
@@ 3.3 Prompt Design
Because the specific examples and their order in the prompt can affect the output of the model, we use exactly the same examples and in the same order across corresponding experiments. To achieve that, we find groups of annotators with significant overlap in the train and in the evaluation sets and use the same samples, including for aggregate and prior. Since the only degree of freedom is the labels, we eliminate many potential confounding factors. We show more details in the Appendix.
@@ 4 Experiments
Figure 3: In-Context Learning performance for annotators and aggregate and their relative improvement compared to the model’s prior as a function of the similarity of each with the model’s prior.
Figure 4: In-Context Learning performance for annotators and their relative improvement compared to the model’s prior as a function of the similarity of each annotator with the aggregate.
Figure 5: Chain-of-Thought performance for annotators and aggregate and their relative improvement compared to the model’s prior as a function of the similarity of each with the model’s prior.
Figure 6: Chain-of-Thought performance for annotators and their relative improvement compared to the model’s prior as a function of the similarity of each annotator with the aggregate.
@@ 4.2 Implementation Details
We use the 4-bit quantized versions of the open-source LLMs through the HuggingFace Wolf et al. (2020) interface for PyTorch. We use LLaMA-2 7B and 70B, LLaMA-3 8B and 70B, GPT-3.5 Turbo, and GPT-4o mini. We chose only models with RLHF Ouyang et al. (2022b) to further reduce confounding factors. We perform 3 runs for each LLM experiment, varying the examples used. Statistical significance is calculated with permutation tests and measured by considering all 3 runs as separate data points. We use random retrieval of examples. We use less shots for CoT given that the reasoning in the prompt increases its length. We generated reasoning chains for each example per annotator and for the aggregate, and use one triplet for CoT. We use one NVIDIA A100 and one V100.
@@ 4.3 Baselines
To establish baseline performance of LLMs compared to smaller, gradient-based methods, we present performance with and without CoT prompting compared to BERT-based Devlin et al. (2018) SpanEmo Alhuzali and Ananiadou (2021). In Figure 1, we demonstrate the significant difference in performance across all the LLMs (45 shots for ICL, 15 shot for CoT) and SpanEmo. In fact, given the very high Jaccard Score and very low F1 scores, results for MFRC indicate close to random performance for the model, even with CoT.
@@ 4.4 Main Results
In all experiments, we present averages on the Y axis, which is (based on) the Micro F1 score for MFRC and the Jaccard Score for GoEmotions. The similarity of the annotators to the prior, as well as the similarity to the aggregate, is measured on the test set, which may differ from the train set.
@@ 6 Limitations
Moreover, datasets with data derived from social media have been criticized as lacking the context for a model—or even humans—to make appropriate judgments about the emotion or morality expressed in them Yang et al. (2023), and techniques to evaluate the correctness of the labels in a dataset have been designed to discard noisy samples Swayamdipta et al. (2020); Mokhberian et al. (2022). Given the subjective nature of the task and the lack of context, such tools could be used to perhaps improve performance. That being said, the interpretation by humans is sufficiently consistent for SpanEmo to achieve better performance than every LLM. We also note that by removing ambiguous examples, we may also remove that which makes these tasks challenging Aroyo and Welty (2015).

### v3

@@ Abstract
In-context Learning (ICL) has become the primary method for performing natural language tasks with Large Language Models (LLMs). The knowledge acquired during pre-training is crucial for this few-shot capability, providing the model with task priors. However, recent studies have shown that ICL predominantly relies on retrieving task priors rather than “learning” to perform tasks. This limitation is particularly evident in complex subjective domains such as emotion and morality, where priors significantly influence posterior predictions. In this work, we examine whether this is the result of the aggregation used in corresponding datasets, where trying to combine low-agreement, disparate annotations might lead to annotation artifacts that create detrimental noise in the prompt. Moreover, we evaluate the posterior bias towards certain annotators by grounding our study in appropriate, quantitative measures of LLM priors. Our results indicate that aggregation is a confounding factor in the modeling of subjective tasks, and advocate focusing on modeling individuals instead. However, aggregation does not explain the entire gap between ICL and the state of the art, meaning other factors in such tasks also account for the observed phenomena. Finally, by rigorously studying annotator-level labels, we find that it is possible for minority annotators to both better align with LLMs and have their perspectives further amplified.111code and data available at: https://github.com/gchochla/aggregation-artifacts-llms
@@ 1 Introduction
While ICL is often contrasted with traditional in-weights learning (i.e., gradient-based updates of the models’ parameters) Kossen et al. (2023); Chan et al. (2022), the ICL abilities of LLMs depend on their general, in-weights prior knowledge, allowing them to perform many tasks in a zero-shot or few-shot manner. Therefore, studying how LLMs leverage the context in relation with their existing knowledge is a prerequisite to understanding ICL.
Prior work found evidence that LLMs may be overly reliant on their prior knowledge, disregarding the demonstrations in the prompt. Specifically, Min et al. (2022) demonstrated that, under certain circumstances, LLMs ignore the provided signal in their prompt in the form of the mapping between inputs and outputs, and instead act as a database of tasks; they focus on the examples and the labels independently to fetch the underlying task Xie et al. (2021): Min et al. (2022) sampled examples and labels independently, and showed very little change in performance. Since no annotations are provided, the setting is given the status of “zero-shot” inference, namely task-recognition zero-shot. While follow-up work Kossen et al. (2023); Wei et al. (2023); Pan et al. (2023); Yoo et al. (2022) has further studied this phenomenon and questioned its generality, more recent work Chochlakis et al. (2024, 2025) has provided further, quantitative evidence for the relative importance of prior and evidence in the posteriors of the models. Specifically, in complex subjective tasks like multilabel emotion or morality recognition, LLMs seem to virtually disregard evidence from the dataset’s mapping in their posterior predictions, even in the form of Chain-of-Thought prompting (CoT; Wei et al. 2022), performing significantly worse than traditional algorithms Alhuzali and Ananiadou (2021); Chochlakis et al. (2023). This may imply an inability to perform annotator modeling, as the same document can receive different but valid annotations by different people.
@@
In this work, we question whether we can efficiently use this crowd truth when modeling these tasks with LLMs. We hypothesize that the aggregation process creates artifacts in the labels provided to the model, and in turn in the prompt, causing the model to ignore the entire input-label mapping as noise. Specifically, aggregation can create inconsistent annotations, since, for instance, different annotators can prevail in different examples, which can cause inconsistencies within or across splits (for a toy example, consult Section C). We create a carefully crafted experimental setting to test our hypothesis with annotator-level labels using ICL and CoT, and try to gauge at the magnitude of this effect. We find that LLMs do indeed tend to consistently favor individual perspectives compared to the aggregate, and in fact favor minority annotators more than majority annotators, who better resemble the aggregate. Our contributions are:
@@ 2.1 In-Context Learning and Priors
ICL Brown et al. (2020) has been used extensively to evaluate LLMs on standard benchmarks Srivastava et al. (2022). It requires no gradient-based interventions, which are otherwise costly to perform for large models, and usually achieves competitive or state-of-the-art performance. The existence of commercial APIs Achiam et al. (2023), coupled with open-weights alternatives Touvron et al. (2023); Dubey et al. (2024) have made ICL an accessible generalist for language tasks and more Liu et al. (2024). Previous work has further studied controlling the reliance on context and in-weights knowledge through the distribution of the training data Chan et al. (2022), examining how to optimally select examples for the prompt Rubin et al. (2022); Gupta et al. (2023), integrating instructions explicitly during training Touvron et al. (2023); Ouyang et al. (2022b), etc. Relevant to our work, prior work has elicited ICL priors by providing random labels for the examples of the prompt Min et al. (2022). The resulting minimal variations in performance indicates that LLMs recognize and retrieve their prior knowledge of the task in the prompt rather than doing any “learning”. Subsequent results challenged the view that LLMs mostly perform task recognition, showcasing a significant degradation in performance when scaling the prompt Kossen et al. (2023) or when focusing on specific tasks instead of aggregates Yoo et al. (2022), and analyzed behavior with certain label manipulations Kossen et al. (2023); Pan et al. (2023); Wei et al. (2023). More recent work, however, suggests that in complex subjective tasks like emotion and morality classification, the prior understanding of the task dominates posterior predictions Chochlakis et al. (2024).
One potential way to augment ICL and overcome the prior bias is with CoT Wei et al. (2022). CoT incorporates the derivation process from input to output explicitly in the prompt, presenting a more human-like reasoning process. This has several advantages, such as making some patterns in the data explicit in the prompt, making model responses more explainable, and potentially directing more computing resources towards more complex problems. However, detailed analysis has cast some doubt on the reliability and the faithfulness of this reasoning technique Lanham et al. (2023); Turpin et al. (2024). Nonetheless, CoT seems to improve the robustness and performance of LLMs across a plethora of tasks, yet follow-up work on subjective tasks showed no improvements and the same bias towards a reasoning prior, especially for larger models Chochlakis et al. (2025). Methods such as Tree of Thoughts Yao et al. (2024) or self-consistency Wang et al. (2022) have experimented with ways to further augment CoT. In this work, we study whether part of the prior bias in both ICL and CoT can be explained by the aggregation process used to derive the labels, and in turn the reasoning process, of the examined benchmarks.
@@ 2.2 Annotator Disagreement and Modeling
Many works have attempted to model individual annotator perspectives instead of the aggregate, like we advocate in this work. For example, researchers used the EM algorithm Dawid and Skene (1979) to assign confidence to each annotator’s evaluations Hovy et al. (2013). Recently, Gordon et al. (2022) concatenated features derived from Transformers Vaswani et al. (2017) with annotator embeddings that incorporate demographics to model individual perspectives. Demographic information has also been incorporated in word embeddings by  Garten et al. (2019). Demographic information and psychological profiles have been statistically examined in text annotations to derive insights into systematic biases Sap et al. (2022). Recent work has tried to filter annotators based on deep learning methods Mokhberian et al. (2022), to model annotators on top of common representations Davani et al. (2022); Mokhberian et al. (2023), and to decrease annotation costs based on agreement Golazizian et al. (2024). Modeling annotators with LLMs has shown limited success, and LLM biases have also been explored Dutta et al. (2023); Abdurahman et al. (2024); Hartmann et al. (2023).
@@ 3 Methodology
We closely follow the methodology and notation of Chochlakis et al. (2025). For a set of examples $\mathcal{X}$, and a set of labels $\mathcal{Y}$, a dataset $\mathcal{D}^{a}$ defines a mapping $f^{a}:\mathcal{X}\rightarrow\mathcal{Y}$, where $a$ denotes a specific annotator or the aggregate, as well as reasoning chains $R^{a}(x)=r,x\in\mathcal{X}$ that explicitly describe $f^{a}$, and therefore $\mathcal{D}^{a}=\{(x,y,r):x\in\mathcal{X},y=f^{a}(x),r=R^{a}(x)\}$, from which we can sample demonstrations with $p(x,y,r)$. We do not differentiate between splits for brevity. Given CoT prompt $S=\{(x_{i},y_{i},r_{i}):(x_{i},y_{i},r_{i})\sim p_{S},i\in[k]\}$ with $k$ demonstrations sampled with distribution $p_{S}$ from $\mathcal{D}^{a}$ without replacement, an LLM produces its own mapping and predictions for the task, denoted as $\hat{f}_{k}(.;p_{S}):\mathcal{X}\rightarrow\mathcal{Y}$. When using regular ICL, we simply drop the reasoning text. For all our experiments, we set the temperature $\tau=0$ to derive deterministic predictions, and vary the examples and/or labels between runs.
@@ 3.1 Similarity and Performance Metrics
To keep evaluations consistent when using API-based LLMs, we rely on similarity measures calculated directly on the final predictions rather than probabilistic measures like logits. Therefore, we use the Jaccard Score (JS), Micro and Macro F1 metrics Mohammad et al. (2018) to evaluate the performance of the models. For consistency, we also use them to quantify the similarity between the predictions from different model runs or annotators, since they are symmetric functions which we apply to interchangeable sets.
Figure 1: Performance comparison between LLMs w/ and w/o Chain-of-Thought prompting compared to Demux (BERT-based) using aggregated labels.
Figure 2: Scatter plot of annotator similarity with aggregate and with prior. Correlation and line fit of data also shown. and bold lines: $p<0.05$.
@@ 3.2 Task Prior Knowledge Proxies via Zero-shot Inference
Here, we precisely define the priors that we use for ICL and CoT. We use the term prior to contrast it with the posterior predictions of the model after evidence (i.e., a specific annotator’s labels) from the dataset have been presented to it. First, we have the true reasoning task-recognition zero-shot222since no annotations are required prior, where the prompt contains $k$ demonstrations sampled with $p^{I}(x,y,r)=p(x)p(y)p(r)$, so text, labels, and reasoning are sampled independently from each other from $\mathcal{D}^{a}$, hence labels and reasoning are irrelevant to the text and each other. This effectively maintains the relationships between labels, which are strong in such multilabel tasks Cowen and Keltner (2017). For ICL, we have the corresponding task-recognition zero-shot prior sampled with $p^{I}(x,y)=p(x)p(y)$, so text and labels are also sampled independently. The similarity of the priors to annotators and aggregate are, therefore, measured by comparing (as described in Section 3.1) the prior predictions $\hat{f}_{k}(.;p_{I})$ with the annotator’s labels, which is equivalent to the prior performance for the annotator, and the posteriors are simply defined as the $\hat{f}_{k}(.;p_{S})$ with the joint sampling distribution $p_{S}$ (meaning we present the gold labels to the model).
@@ 3.3 Prompt Design
Because the specific examples and their order in the prompt can affect the output of the model, we use exactly the same examples and in the same order across corresponding experiments. To achieve that, we find groups of annotators with significant overlap in the train and in the evaluation sets and use the same samples, including for aggregate and prior. Since the only degree of freedom is the labels, we eliminate the aforementioned confounding factors, among others (more details in Section D).
@@ 4 Experiments
Figure 3: In-Context Learning performance for annotators and aggregate and their relative improvement compared to the model’s prior as a function of the similarity of each with the model’s prior. Correlation and line fit of data also shown. and bold lines: $p<0.05$.
Figure 4: In-Context Learning performance for annotators and their relative improvement compared to the model’s prior as a function of the similarity of each annotator with the aggregate. Correlation and line fit of data also shown. and bold lines: $p<0.05$.
Figure 5: Chain-of-Thought performance for annotators and aggregate and their relative improvement compared to the model’s prior as a function of the similarity of each with the model’s prior. Correlation and line fit of data also shown. and bold lines: $p<0.05$.
Figure 6: Chain-of-Thought performance for annotators and their relative improvement compared to the model’s prior as a function of the similarity of each annotator with the aggregate. Correlation and line fit of data also shown. and bold lines: $p<0.05$.
@@ 4.2 Implementation Details
We use the 4-bit quantized versions of the open-source LLMs through the HuggingFace Wolf et al. (2020) interface for PyTorch. We use LLaMA-2 7B and 70B, LLaMA-3 8B and 70B, GPT-3.5 Turbo, and GPT-4o mini. We chose only models with RLHF Ouyang et al. (2022b) for uniformity. We perform 3 runs for each LLM experiment, varying the examples used. Statistical significance is calculated with permutation tests and measured by considering all 3 runs as separate data points. We use random retrieval of examples. We use less shots for CoT given the increases in length in the prompt. We generated reasonings for each example per annotator and for the aggregate. For details on our CoT annotations, see Section B in the Appendix. We use one NVIDIA A100 and one V100.
@@ 4.3 Baselines
To establish baseline performance of LLMs compared to smaller, gradient-based methods, we present performance with and without CoT prompting compared to BERT-based Devlin et al. (2018) Demux Chochlakis et al. (2023). In Figure 1, we demonstrate the significant difference in performance across all the LLMs (45 shots for ICL, 15 shot for CoT) and Demux. In fact, given the very high JS and very low F1 scores, results for MFRC indicate close to random performance for the model, so we choose to use Micro F1 for MFRC, as opposed to JS for GoEmotions.
@@ 4.4 Main Results
In all experiments, we present averages. The similarity of the annotators to the prior, as well as the similarity to the aggregate, is measured on the test set, which may differ from the train set.
@@ 6 Limitations
A potential confounding factor that we do not control for is the quantization, as previous work has reported significant decreases in performance from it Marchisio et al. (2024). We note, first, that there is no reason for the quantization to affect our results in a nonuniform way, second, that we perform the quantization because of obvious computational constraints, and third, it is possible that even some API-based models are served quantized (e.g., Turbo versions). For these reasons, we believe that quantized performance is representative of LLM performance in realistic scenarios. Moreover, this work does not aim to establish the performance of LLMs in these subjective tasks compared to other models, but rather to compare within the LLMs themselves. Nonetheless, we perform experiments with various levels of quantization of LLaMA-3 8B, and show the results in Table 3 in the Appendix.
@@
Moreover, datasets with data derived from social media have been criticized as lacking the context for a model—or even humans—to make appropriate judgments about the emotion or morality expressed in them Yang et al. (2023), and techniques to evaluate the correctness of the labels in a dataset have been designed to discard noisy samples Swayamdipta et al. (2020); Mokhberian et al. (2022). Given the subjective nature of the task and the lack of context, such tools could be used to perhaps improve performance. That being said, the interpretation by humans is sufficiently consistent for Demux to achieve better performance than every LLM. We also note that by removing ambiguous examples, we may also remove that which makes these tasks challenging Aroyo and Welty (2015).